2. Bob & Sue Realty (BSR) is a publicly traded REIT that has no debt and a current dividend yield of 8%, with a current share price/earnings multiple of 12.5. The current consensus expectation among stock analysts who follow BSR is that BSR can provide a long-term average growth rate in its dividends per share of 5% per year.

a. What is BSR’s plowback ratio (i.e., what proportion of its earnings does it retain and not pay as dividends)? 12 points.

b. Assuming the stock market agrees with these analysts’ expectations, what is BSR’s firm-level average cost of capital? [Hint: As BSR has no debt, you can use the GGM directly to answer this question.] 12 points.
Answer #1

a.Price/Earning =P/E=12

Price =P=12*Earning

Assume ,

Dividend payout ratio=r

Amount of dividend per share =r*E

Dividend Yield =Dividend/Price =(r*E/P)=0.08

r=0.08*(P/E)=0.08*12=0.96

Dividend payout ratio=0.96

Plowback ratio=1-dividend payout ratio=1-0.96=0.04

b.

Dividend Yield =Dividend/Price =0.08

Growth rate =g=5%=0.05

Required Return =R

P=Dividend/(R-g)

R=(Dividend/Price)+g=Dividend Yield+g

Cost of Capital =Required Return=R=0.08+0.05=0.13

Average Cost of Capital =13%